# Multi-feature Reconstruction Network using Crossed-mask Restoration for Unsupervised Anomaly Detection

### v1

@@ Abstract
Unsupervised anomaly detection using only normal samples is of great significance for quality inspection in industrial manufacturing. Although existing reconstruction-based methods have achieved promising results, they still face two problems: poor distinguishable information in image reconstruction and well abnormal regeneration caused by model over-generalization ability. To overcome the above issues, we convert the image reconstruction into a combination of parallel feature restorations and propose a multi-feature reconstruction network, MFRNet, using crossed-mask restoration in this paper. Specifically, a multi-scale feature aggregator is first developed to generate more discriminative hierarchical representations of the input images from a pre-trained model. Subsequently, a crossed-mask generator is adopted to randomly cover the extracted feature map, followed by a restoration network based on the transformer structure for high-quality repair of the missing regions. Finally, a hybrid loss is equipped to guide model training and anomaly estimation, which gives consideration to both the pixel and structural similarity. Extensive experiments show that our method is highly competitive with or significantly outperforms other state-of-the-arts on four public available datasets and one self-made dataset.
@@
Anomaly detection aims to identify and locate anomalous regions in images [1], which is a fundamental yet challenging task in visual understanding that has a wide range of real-world applications, such as product quality monitoring [2, 3, 4], medical health diagnosis [5, 6], video surveillance [7, 8, 9], and so on. In the past decades, machine vision-based methods [10, 11] have gained significant attention, and are increasingly replacing the traditional manual inspection, which is time-consuming, subjective and inaccurate. However, these previous works generally rely on the extraction of handcrafted features based on the characteristics of a specific situation, limiting the generalization ability of the algorithms. Since the unprecedented breakthrough of deep learning in computer vision, great efforts have been made to apply it to high-quality anomaly detection. Nevertheless, in most real application scenarios, it is easy to acquire normal examples, but unrealistic to collect and label all types of abnormality, which makes it impossible to apply well-established full-supervised learning methods [12, 13]. Therefore, unsupervised learning methods that only require normal data for training are attracting more and more attention. The two classical paradigms for unsupervised anomaly detection methods are feature-based and reconstruction-based.
@@
Secondly, the excessive generalization ability of the model makes it challenging to ensure a poor reconstruction of abnormalities. More specifically, reconstruction-based methods assume that reconstruction only succeeds when samples are normal. But due to the strong generalization ability of neural networks, anomalies could be regenerated well unexpectedly in practice, thus resulting in false detection. Several empirical methodologies [7, 17] have been used to address this problem, but the effectiveness remains limited. Consequently, exploring an effective reconstruction model to prevent anomalies from appearing in the reconstructed images remains an open question.
@@
2. We consider the reconstruction task as a restoration problem to break information equivalence, which can learn more distinguishable semantic information and thus reduce over-generalization to anomalies.
@@ II-A Feature-based approaches
For the feature extraction modules, previous studies [10, 11] utilized a number of simple handcrafted feature descriptors, such as texture features or edge features, to represent the images. However, they generally fail to yield impressive detection performance, due to the lack of semantically discriminative features. With the rapid development of deep learning, in recent years, the trend has been shifted to use deep learning methods as feature extractors. SPADE [18] could be the first work that directly used a deep neural network pre-trained on the large-scale image database ImageNet [19] to obtain better feature representation. Subsequently, more and more researchers [14, 20, 21, 22, 23] began to experiment with a variety of popular pre-trained networks, e.g., ResNet, Wide-ResNet or EfficientNet, and also produced decent anomaly detection results. Most of these studies discussed the effect of using different layer features on the experimental results, and declared that the ensemble of multiple feature representations of the networks can offer some benefits. However, there is still no general consensus on how to choose the appropriate feature layers for aggregation to cope with different scenarios. Besides using pre-trained deep networks, significant efforts had also been invested in designing proper self-supervised learning tasks to learn meaningful representations from scratch. The popular proxy tasks include predicting the relative position of a pair of random image patches [24, 25], solving jigsaw puzzles of the randomly disrupted image patches [26], restoring the erased attributes of raw image (color and orientation information) [16], binary classification between the normal samples and the artificial pseudo-anomaly samples [15, 27], obtaining the learned latent features from deep auto-encoders via CNN [28, 29] or vision transformer [30], and so on. Although these elaborate auxiliary tasks have shown great potential in extracting semantic features, the learned representation capacity will be greatly limited by the limited normal training data. Furthermore, such methods require an additional training phase, which is not off-the-shelf like pre-trained features, and they also don’t exhibit significant advantages over the ones using pre-trained features [23]. Very recent works [32, 33] starting with [31] are able to further transform the pre-trained features of normal data into a more simple well-defined distribution, typically Gaussian distribution, by using normalizing flow-based models. Due to being projected to a more tractable distribution, they generally showed favorable detection performance.
@@ II-B Reconstruction-based approaches
The most fundamental architecture of reconstruction-based methods, auto-encoder, is essentially an encoder-decoder network. The encoder manages to learn meaningful representations of normal inputs into a latent space, and the decoder attempts to regenerate the inputs from this low-dimensional space. The abnormal data unseen during the training are expected to not be reconstructed as accurately as normal ones, hence making higher reconstruction errors. Over the past few years, various variants of auto-encoder networks had been extensively explored for anomaly detection task. Chow et al. [37] proposed to adopt convolutional AE for anomaly recognition, in which the hidden layers are implement by convolution. While Chen et al. [38] implemented auto-encoder by the emerging Transformer structure to capture more non-local feature information. Further study [39] attempted to inspect anomaly by the variational AE, but did not show significant performance gains over the convolutional AE. Additionally, the training of variational AE is relatively unstable. To improve the robustness of convolutional AE, Mei et al. [40] introduced noise interference into the input images for model training and obtained a better detection performance. Although these models are supposed to only reconstruct normal images accurately, the powerful generalization capabilities of neural networks sometimes result in anomalies being reproduced well, thus causing misdetection. Lots of efforts have been devoted to tackling this excessive generalization issue. [7] believed that skip connections in most auto-encoders may simply copy the information of abnormal images, and thus proposed a U2-net-like backbone network without skip connections. Some works [17, 41] introduced a memory module into auto-encoder to record the normal patterns of the training samples. Given a defective input, the proposed memory-augmented AE will try to reproduce images using normal features stored in the memory module, reducing the possibility of recreating anomalies. To restrict the representation learning in the latent space explicitly, Liu et al. [4] and Niu et al. [42] proposed the novel auto-encoder with dual prototype loss, which were able to encourage the latent representation of encoder to keep closer to their own prototype, i.e., the center of the latent representation of training images. Very recently, Fei et al. [16] held a view that the cause of anomaly reconstruction may be the failure to extract semantic features in the vanilla AE models. They suggested that the AE models tend to simply compress images instead of learning the abstract semantic representation, due to the information equivalence, namely the same information between the input and output. Later on, many following researchers [43, 44, 45] adopted image inpainting for anomaly localization to break the information equivalence and thus obtained more semantic understanding of the image. In these methods, parts of the sample images are covered by pre-specified masks, and then these hide regions are restored conditioned on their surroundings by the inpainting models. During the testing, anomaly scores are determined by the discrepancy between the masked regions and the corresponding repairs. This effective paradigm also inspires our work to formulate the anomaly detection task as a restoration problem through cross-masking and then recovering. The vast majority of reconstruction-based approaches just perform reconstruction in raw image space, which is insufficient for complex scenarios. To introduce more image information into the reconstruction process, some attempts have been made to execute reconstruction in the space of pre-trained features [38, 46], image pyramids [40], multi-frequency [47, 48] or texture–structure features [49]. Meanwhile, there are some works trying to leverage different distance metrics to assess the difference between inputs and reconstructed outputs. A simple or loss was used in [37, 38, 46] for model training, however, this ignored the inter-dependencies between the neighboring regions. [43, 45] and [50] introduced structural similarity losses to give a better measure for image similarity.
@@ III-B Crossed-mask restoration network
After the multi-scale features of the image are ready, we leverage a crossed-mask restoration network to extend the reconstruction-based approach for anomaly localization. Specifically, a set of complementary crossed masks are generated to formulate a restoration task, followed by a restoration network trained to predict the covered regions. In this way, the difference between the masking region and the corresponding restoration is going to be significant for anomalous regions. By covering and restoring the partial regions of the input, our method can gain a much deeper semantic understanding of the scene, thus limiting the over-generalization capability to prevent anomalies from being reconstructed well.
@@ IV-A2 Datasets description
e) Fabric-US dataset: This is a novel dataset of fabric images created by ourselves, which is available to the public 111https://pan.baidu.com/s/19gZoBxI9AUCQUTTBxyq8mA?pwd=0927. It is the unsupervised version of previous Fabric dataset [58], thus named Fabric-US (UnSupervised). Within our Fabric-US dataset, 180 normal images are newly added to serve as the training set, while 400 typical defect images in the original dataset are adopted as abnormal samples of the testing set. The resolution of each raw image is 512512 pixels. Since the weak defect features, this dataset becomes challenging.
@@ IV-C1 Architecture ablation
As listed in Table V, the baseline holds an unsatisfactory result and leaves room for improvement, due to the limited distinguishable information in image space and overly strong generalization ability of the model. Then, instead of reconstructing the raw image as in baseline, we attempt to reconstruct the corresponding feature map, which can provide more distinguishable representation. From the second row of Table V, we can observe that this modification is extremely beneficial, achieving a reduction of 10.7 in MAE, and contributing gains of 9.3, 7.6 and 37.5 in AUROC, ACC and F1 metrics respectively, compared to the baseline. We can also see that the reconstruction of the feature map can significantly suppress erroneously reported anomalies in Fig. 5(d). Next, the reconstruction problem is converted into a restoration problem by introducing CG module, and the convolutional network in baseline is still utilized to repair the masked feature map. We observe a substantial improvement in AUROC (86.68 v.s. 91.79), MAE (7.93 v.s. 7.65), ACC (85.73 v.s. 88.56) and F1 (42.85 v.s. 50.79). It can also be noted from Fig. 5(e) that the problem of anomalies being well reconstructed has been effectively addressed. Besides, when using the proposed restoration network with hybrid transformer blocks, the overall performance is further increased, showing it enjoys the benefits of both long-range contextual information and low-level details. A similar observation can be found in Fig. 5(f).

### v2

# Multi-feature Reconstruction Network using Crossed-mask Restoration for Unsupervised Industrial Anomaly Detection

@@ Abstract
Unsupervised anomaly detection using only normal samples is of great significance for quality inspection in industrial manufacturing. Although existing reconstruction-based methods have achieved promising results, they still face two problems: poor distinguishable information in image reconstruction and well abnormal regeneration caused by model under-regularization. To overcome the above issues, we convert the image reconstruction into a combination of parallel feature restorations and propose a multi-feature reconstruction network, MFRNet, using crossed-mask restoration in this paper. Specifically, a multi-scale feature aggregator is first developed to generate more discriminative hierarchical representations of the input images from a pre-trained model. Subsequently, a crossed-mask generator is adopted to randomly cover the extracted feature map, followed by a restoration network based on the transformer structure for high-quality repair of the missing regions. Finally, a hybrid loss is equipped to guide model training and anomaly estimation, which gives consideration to both the pixel and structural similarity. Extensive experiments show that our method is highly competitive with or significantly outperforms other state-of-the-arts on four public available datasets and one self-made dataset.
@@
Anomaly detection aims to identify and locate anomalous regions in images [1], which is a fundamental yet challenging task in visual understanding that has a wide range of real-world applications, such as product quality monitoring [2, 3, 4], medical health diagnosis [5, 6], video surveillance [7, 8], and so on. In the past decades, machine vision-based methods [9, 10] have gained significant attention, and are increasingly replacing the traditional manual inspection, which is time-consuming, subjective and inaccurate. However, these previous works generally rely on the extraction of handcrafted features based on the characteristics of a specific situation, limiting the generalization ability of the algorithms. Since the unprecedented breakthrough of deep learning in computer vision, great efforts have been made to apply it to high-quality anomaly detection. Nevertheless, in most real application scenarios, it is easy to acquire normal examples, but unrealistic to collect and label all types of abnormality, which makes it impossible to apply well-established full-supervised learning methods [11, 12]. Therefore, unsupervised learning methods that only require normal data for training are attracting more and more attention. The two classical paradigms for unsupervised anomaly detection methods are feature-based and reconstruction-based.
@@
Secondly, the under-regularization ability of the reconstruction model makes it challenging to ensure a poor reconstruction of abnormalities. More specifically, reconstruction-based methods assume that reconstruction only succeeds when samples are normal. But due to the strong generalization ability of neural networks, anomalies could be regenerated well unexpectedly in practice, thus resulting in false detection. Several empirical methodologies [7, 16] have been used to address this problem, but the effectiveness remains limited. Consequently, exploring an effective regularization method to prevent the reconstruction model from regenerating anomalies remains an open question.
@@
2. We consider the reconstruction task as a restoration problem to break information equivalence, which can learn more distinguishable semantic information and regularize the reconstruction model to prevent abnormal reconstruction.
@@ II-A Feature-based approaches
For the feature extraction modules, previous studies [9, 10] utilized a number of simple handcrafted feature descriptors, such as texture features or edge features, to represent the images. However, they generally fail to yield impressive detection performance, due to the lack of semantically discriminative features. With the rapid development of deep learning, in recent years, the trend has been shifted to use deep learning methods as feature extractors. SPADE [17] could be the first work that directly used a deep neural network pre-trained on the large-scale image database ImageNet [18] to obtain better feature representation. Subsequently, more and more researchers [13, 19, 20, 21, 22] began to experiment with a variety of popular pre-trained networks, e.g., ResNet, Wide-ResNet or EfficientNet, and also produced decent anomaly detection results. Most of these studies discussed the effect of using different layer features on the experimental results, and declared that the ensemble of multiple feature representations of the networks can offer some benefits. However, there is still no general consensus on how to choose the appropriate feature layers for aggregation to cope with different scenarios. Besides using pre-trained deep networks, significant efforts had also been invested in designing proper self-supervised learning tasks to learn meaningful representations from scratch. The popular proxy tasks include predicting the relative position of a pair of random image patches [23, 24], solving jigsaw puzzles of the randomly disrupted image patches [25], restoring the erased attributes of raw image (color and orientation information) [15], binary classification between the normal samples and the artificial pseudo-anomaly samples [14, 26], obtaining the learned latent features from deep auto-encoders via CNN [27, 28] or vision transformer [29], and so on. Although these elaborate auxiliary tasks have shown great potential in extracting semantic features, the learned representation capacity will be greatly limited by the limited normal training data. Furthermore, such methods require an additional training phase, which is not off-the-shelf like pre-trained features, and they also don’t exhibit significant advantages over the ones using pre-trained features. Very recent works [31, 32] starting with [30] are able to further transform the pre-trained features of normal data into a more simple well-defined distribution, typically Gaussian distribution, by using normalizing flow-based models. Due to being projected to a more tractable distribution, they generally showed favorable detection performance.
@@ II-B Reconstruction-based approaches
The most fundamental architecture of reconstruction-based methods, auto-encoder, is essentially an encoder-decoder network. The encoder manages to learn meaningful representations of normal inputs into a latent space, and the decoder attempts to regenerate the inputs from this low-dimensional space. The abnormal data unseen during the training are expected to not be reconstructed as accurately as normal ones, hence making higher reconstruction errors. Over the past few years, various variants of auto-encoder networks had been extensively explored for anomaly detection task. Chow et al. [36] proposed to adopt convolutional AE for anomaly recognition, in which the hidden layers are implement by convolution. While Chen et al. [37] implemented auto-encoder by the emerging Transformer structure to capture more non-local feature information. Further study [38] attempted to inspect anomaly by the variational AE, but did not show significant performance gains over the convolutional AE. Additionally, the training of variational AE is relatively unstable. To improve the robustness of convolutional AE, Mei et al. [39] introduced noise interference into the input images for model training and obtained a better detection performance. Although these models are supposed to only reconstruct normal images accurately, the under-regularization of neural networks sometimes result in anomalies being reproduced well, thus causing misdetection. Lots of efforts have been devoted to tackling this excessive generalization issue. [7] believed that skip connections in most auto-encoders may simply copy the information of abnormal images, and thus proposed a U2-net-like backbone network without skip connections. Some works [16, 40] introduced a memory module into auto-encoder to record the normal patterns of the training samples. Given a defective input, the proposed memory-augmented AE will try to reproduce images using normal features stored in the memory module, reducing the possibility of recreating anomalies. To restrict the representation learning in the latent space explicitly, Liu et al. [4] and Niu et al. [41] proposed the novel auto-encoder with dual prototype loss, which were able to encourage the latent representation of encoder to keep closer to their own prototype, i.e., the center of the latent representation of training images. Very recently, Fei et al. [15] held a view that the cause of anomaly reconstruction may be the failure to extract semantic features in the vanilla AE models. They suggested that the AE models tend to simply compress images instead of learning the abstract semantic representation, due to the information equivalence, namely the same information between the input and output. Later on, many following researchers [42, 43, 44] adopted image inpainting for anomaly localization to break the information equivalence and thus obtained more semantic understanding of the image. In these methods, parts of the sample images are covered by pre-specified masks, and then these hide regions are restored conditioned on their surroundings by the inpainting models. During the testing, anomaly scores are determined by the discrepancy between the masked regions and the corresponding repairs. This effective paradigm also inspires our work to formulate the anomaly detection task as a restoration problem through cross-masking and then recovering. The vast majority of reconstruction-based approaches just perform reconstruction in raw image space, which is insufficient for complex scenarios. To introduce more image information into the reconstruction process, some attempts have been made to execute reconstruction in the space of pre-trained features [37, 45], image pyramids [39], multi-frequency [46, 47] or texture–structure features [48]. Meanwhile, there are some works trying to leverage different distance metrics to assess the difference between inputs and reconstructed outputs. A simple or loss was used in [36, 37, 45] for model training, however, this ignored the inter-dependencies between the neighboring regions. [42, 44] and [49] introduced structural similarity losses to give a better measure for image similarity.
@@ III-B Crossed-mask restoration network
After the multi-scale features of the image are ready, we leverage a crossed-mask restoration network to extend the reconstruction-based approach for anomaly localization. Specifically, a set of complementary crossed masks are generated to formulate a restoration task, followed by a restoration network trained to predict the covered regions. In this way, the difference between the masking region and the corresponding restoration is going to be significant for anomalous regions. By covering and restoring the partial regions of the input, our method can gain a much deeper semantic understanding of the scene, and regularizes the reconstruction model to prevent anomalies from being reconstructed well.
@@ IV-A2 Datasets description
e) Fabric-US dataset: This is a novel dataset of fabric images collected by our proposed acquisition equipment shown in Fig.5. The acquisition system primarily consists of line-scan digital cameras, strip lighting, and transmission roller. The line-scan cameras are equipped with high-precision CCD sensors, offering a lateral resolution of 4K and capable of achieving a maximum scanning speed of 16 meters per second. The strip light is strategically positioned perpendicular to the fabric surface, ensuring sufficient brightness to enhance the visibility of defects and make them more discernible. Due to the inherent softness and flexibility of fabric, the transmission roller facilitates the controlled movement and rolling of the material during the scanning process, ensuring stability and consistent image capture. This integrated system provides a reliable and efficient platform for high-quality fabric image acquisition, supporting detailed analysis and defect detection. Finally, 180 normal images are captured to serve as a training set, and 400 typical defect images are adopted as abnormal samples for the testing. The resolution of each raw image is 512512 pixels. Since the weak defect features, this dataset becomes challenging. This dataset is the unsupervised version of previous Fabric dataset [57], thus named Fabric-US (UnSupervised), and is available to the public111https://pan.baidu.com/s/19gZoBxI9AUCQUTTBxyq8mA?pwd=0927.
@@ IV-C1 Architecture ablation
As listed in Table V, the baseline holds an unsatisfactory result and leaves room for improvement, due to the limited distinguishable information in image space and under-regularization of the model. Then, instead of reconstructing the raw image as in baseline, we attempt to reconstruct the corresponding feature map, which can provide more distinguishable representation. From the second row of Table V, we can observe that this modification is extremely beneficial, achieving a reduction of 10.7 in MAE, and contributing gains of 9.3, 7.6 and 37.5 in AUROC, ACC and F1 metrics respectively, compared to the baseline. We can also see that the reconstruction of the feature map can significantly suppress erroneously reported anomalies in Fig. 6(d). Next, the reconstruction problem is converted into a restoration problem by introducing CG module, and the convolutional network in baseline is still utilized to repair the masked feature map. We observe a substantial improvement in AUROC (86.68 v.s. 91.79), MAE (7.93 v.s. 7.65), ACC (85.73 v.s. 88.56) and F1 (42.85 v.s. 50.79). It can also be noted from Fig. 6(e) that the problem of anomalies being well reconstructed has been effectively addressed. Besides, when using the proposed restoration network with hybrid transformer blocks, the overall performance is further increased, showing it enjoys the benefits of both long-range contextual information and low-level details. A similar observation can be found in Fig. 6(f).